the length of a rectangle is 2 feet less than three times the width. the area of the rectangle is 40 square feet. find the length and the width of the rectangle.
Answer by Cromlix(4381)

Hi there,
Make width = 'x'
Length therefore, = 3x - 2
Area = Width * Length (* Times)
40 = x(3x - 2)
40 = 3x^2 - 2x
3x^2 - 2x - 40 = 0
Factorise
(3x + 10)(x - 4) = 0
3x + 10 = 0
3x = -10
x = -10/3 (Disregard as -ve)
x - 4 = 0
x = 4
Width = 4 feet
Length = 10 feet.
Hope this helps :-)